OpenStudy (anonymous):

what is 8x = 4y =16 step by step in slope intercept form

OpenStudy (anonymous):

its y = mx + b but you could also do y=x2-x1+m+y2-y1

OpenStudy (anonymous):

Its been a while ive done this but I think I mixed the ys and xs up I think you do y=y2-y1+m(slope)+x2-x1